Chris got 48 out of 55 in a test and Kim got 33 out of 38 in a different test. The students wanted to know which student would achieve a higher percentage based on these results. Which of the following is true?

a. They would get the same percentage

b. Kim would get a higher percentage than Chris

c. Chris would get a higher percentage than Kim

d. It is not possible to tell
Answer #1

Chris percentage in exam = (48/55) x 100 = 87.27%

Kim percentage in exam = (33/38) x 100 = 86.84%

Ans. C. Chris would get a higher percentage than kim